OpenStudy (anonymous):

Two equations are given below: a − 3b = 4 a = b − 2 What is the solution to the set of equations in the form (a, b)? (−2, −2) (−3, −1) (−9, −7) (−5, −3)

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

can anyone help?

OpenStudy (anonymous):

since you know what a is equal to plug its value into the first equation so that you have the first equation only in terms of b, then solve for b

OpenStudy (anonymous):

@AliSurfCat ok?

OpenStudy (anonymous):

I really dont know

OpenStudy (anonymous):

what is a equal to? according to the second equation

OpenStudy (anonymous):

thats the thing I dont know:(

OpenStudy (anonymous):

if i show you this, z = 10, what is z equal to?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

perfect, now if i show you this, a = b − 2, what is a equal to?

OpenStudy (anonymous):

a equals 4

OpenStudy (anonymous):

a equals b - 2

OpenStudy (anonymous):

and be equals 6

OpenStudy (anonymous):

no, a equals b - 2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

ok lets make sure you understand this, if i show you j = 5i +z - 2, what is j equal to?

OpenStudy (anonymous):

5i +z - 2

OpenStudy (anonymous):

perfect :)

OpenStudy (anonymous):

ok back to your question, so we know a is equal to b - 2

OpenStudy (anonymous):

since a is equal to b -2 can you please replace a in the first equation with b -2 since they equal the same thing

OpenStudy (anonymous):

ok, b-2 − 3b = 4

OpenStudy (anonymous):

perfect :)

OpenStudy (anonymous):

now can you solve for b? with that equation b -2 -3b = 4

OpenStudy (anonymous):

cna you help me on this one?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

to solve for b we need to find out what b equals, so we have an equation that has only numbers and the variable b, so we need to perform some arithmetic on the equation to get b by itself on one side of the equals sign

OpenStudy (anonymous):

so lets start by only getting terms with a b in them on one side, so if you have b - 2 - 3b = 4, what would be a good step to take so that there are only terms with b in them on the left side?

OpenStudy (anonymous):

how can i get rid of the -2 on the left side?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what happens if we add 2 on both sides of the equation, what do you get?

OpenStudy (anonymous):

you get 6

OpenStudy (anonymous):

i get 6 on the right hand side of the equals sign but what about on the left side?

OpenStudy (anonymous):

b - 2 + 2 - 3b = 4 + 2 b - 2 + 2 - 3b = 6

OpenStudy (anonymous):

hey I have a few minutes left to turn this in please help

OpenStudy (anonymous):

i am helping, if you just want the answer, i cant do that

OpenStudy (anonymous):

I dont but think it is A

OpenStudy (anonymous):

good luck bud

OpenStudy (anonymous):

ok well thanks

OpenStudy (anonymous):

if you change your mind and want to keep working through it let me know

OpenStudy (anonymous):

I do :)

OpenStudy (anonymous):

ok so we added 2 to both sides to get this b - 2 + 2 - 3b = 4 + 2 b - 2 + 2 - 3b = 6 can you simplify this and tell me what we get?

OpenStudy (anonymous):

b-4-3b=6

OpenStudy (anonymous):

close, but -2 + 2 is not equal to -4

OpenStudy (anonymous):

ohhh sorry got a little confused

OpenStudy (anonymous):

ok so -2 + 2 = what?

OpenStudy (anonymous):

it eaquals 0

OpenStudy (anonymous):

perfect, now we have b + 0 - 3b=6 which equals b - 3b = 6

OpenStudy (anonymous):

can you simplify b - 3b = 6 by combining the b terms?

OpenStudy (anonymous):

how do I do that

OpenStudy (anonymous):

if i have 1b and i take away 3b's how many b's do i have?

OpenStudy (anonymous):

do I just do 3b=6???

OpenStudy (anonymous):

1b - 3b

OpenStudy (anonymous):

--2b

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so we are almost done we have -2b = 6, now how can you get b by itself?

OpenStudy (anonymous):

how can we get rid of the -2?

OpenStudy (anonymous):

turn it into a 1

OpenStudy (anonymous):

by adding 2 on both sides

OpenStudy (anonymous):

by dividing both sides by -2

OpenStudy (anonymous):

yes thats what I meant lol

OpenStudy (anonymous):

-2/-2 = 1 so (-2b)/-2 = 6/(-2) 1b = 6/(-2) b = 6 / (-2)

OpenStudy (anonymous):

so what is 6 / (-2), 6 divided by -2

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

perfect, so b = -3, so now you know the answer, do you see an answer where the second number is -3? (a, b) (a, -3)

OpenStudy (anonymous):

which one is it?

OpenStudy (anonymous):

we can find a very easily also looking at the second equation we have a = b - 2 so if i plug in b i get a = -3 -2

OpenStudy (anonymous):

so it is D?

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

yay thanks so much

OpenStudy (anonymous):

you're welcome, please do some practice on this stuff or your future classes in college will be very difficult, visit khan academy and get help on here and other sites also

OpenStudy (anonymous):

I do

OpenStudy (anonymous):

good :)

OpenStudy (anonymous):

have a good day

OpenStudy (anonymous):

you too

OpenStudy (theviper):

nice to see too much long conversation in such a simple question